# Linear-Size Ancilla Systems for Logical Measurements in QLDPC Codes

### v1

@@ 3.5 Joint $X$ Measurement
With the help of this lemma, we can show that the merged code indeed has $k-1$ logical qubits.
@@
In general, it would be nice if given individual ancilla systems for measuring $\bar{X}_{1}$ and $\bar{X}_{2}$ with only the guarantee that they have fault distance at least $d$, they can be connected via a bridge to measure $\bar{X}_{1}\bar{X}_{2}$ with fault distance at least $d$. Here $\bar{X}_{1}$ and $\bar{X}_{2}$ could be in the same codeblock or different codeblocks. Unfortunately, if they are in the same codeblock, there may exist another logical $\bar{X}_{3}$ (inequivalent to $\bar{X}_{1}$, $\bar{X}_{2}$, and $\bar{X}_{1}\bar{X}_{2}$) with weight larger than $d$ whose weight is reduced to exactly $d$ by both the $\bar{X}_{1}$ and $\bar{X}_{2}$ systems individually. In that case, the joint system could reduce the weight of $\bar{X}_{3}$ to below $d$ by simply multiplying by the same checks from the individual systems that reduced the weight of $\bar{X}_{3}$, provided there is sufficient cancellation of support on the bridge. Thus, the merged code $\mathcal{G}_{XX}$ does not necessarily have good code distance. This suggests we can only prove something very general (without assuming more structure or expansion in the Tanner graph) in the case in which $\bar{X}_{1}$ and $\bar{X}_{2}$ are in separate codeblocks.
Another problem is that in the merged code for measuring $\bar{X}_{1}\bar{X}_{2}$, both $\bar{X}_{1}$ and $\bar{X}_{2}$ are still logical operators. However, it is not possible to use only the good fault distance properties of the individual ancilla systems to show the weights of $\bar{X}_{1}$ and $\bar{X}_{2}$ cannot be reduced since they are stabilizers in their respective individual systems.
We prove the following theorem about the fault distances of the joint ancilla system with a combination of assumptions we feel are reasonable. There are probably other sets of assumptions, perhaps including expansion, that could be turned into similar theorems. In specific and sufficiently small cases, it is possible to just compute the distances of the joint system, as we do in Section 4.
@@
Let $\mathcal{S}$ be the stabilizer group of the original codes and each ancilla qubit, including bridge qubits, in |$\ket{0}$, and $\mathcal{S}_{XX}$ be the stabilizer group of the joint merged code, Figure 6. Also let $\mathcal{L}$ be the group of logical operators of the unmeasured $k-1$ qubit logical space and =$\mathcal{L}^{*}=\mathcal{L}\setminus\{I\}$. The versions of Lemmas 9 and 10 that apply to the joint code imply that the measurement fault distance is $\min(R,d_{Z}(\bar{Z}_{1}\mathcal{L}\mathcal{S}))$ and the logical fault distance is $\min(d_{Z}(\mathcal{L}^{*}\mathcal{S}),d_{X}(\mathcal{L}^{*}\mathcal{S}_{XX}))$. The original code $\mathcal{S}$ has code distance at least $d$ – if it did not, then the individual systems would not have fault distances at least $d$ – and so $d_{Z}(\bar{Z}_{1}\mathcal{L}\mathcal{S})\geq d$ and $d_{Z}(\mathcal{L}^{*}\mathcal{S})\geq d$. To complete the proof, it remains only to show that the $X$-distance of the joint merged code is good, i.e. $d_{X}(\mathcal{L}^{*}\mathcal{S}_{XX})\geq d$.
Since $|X(B)S_{1}|$ is at least the weight of factors in line 33, we proceed by ignoring the factors in line 34.
3.6 Gauged $Y$ Ancilla System
@@
$G=\left(\begin{array}[]{ccc}f^{Z\top}&f^{X\top}&0\\ F^{Z\top}&0&S^{Z}\\ 0&F^{X\top}&S^{X}\end{array}\right).$ (35)
@@
We claim that solutions $(v_{X},v_{Z},0)$ with $v_{B}=0$ are the sum of solutions in $U^{X}$ and $U^{Z}$. Note that, for any $v_{X}$, applying $X(v_{X}\in C_{1}^{X})$ flips an even number of checks from $V_{1}^{Z}\cup q_{1}$ (see Lemma 8). Therefore, because $F^{Z\top}v_{X}^{\top}=0$, meaning $v_{X}$ flips no checks in $V_{1}^{Z}$, $v_{X}$ also does not flip $q_{1}$ or $f^{Z\top}v_{X}^{\top}=0$. This shows $(v_{X},0,0)$ is a solution to Eq. (36). Likewise, it can be argued $(0,v_{Z},0)$ is a solution.
@@
Using Lemma 8, we can find a vector $v_{X}$ such that $f^{Z\top}v_{X}^{\top}=1$ and $F^{Z\top}v_{X}^{\top}=\hat{e}_{i_{X}}^{\top}$ because $X(v_{X}\in C_{1}^{X})$ flips exactly two checks, $i_{X}$ and $q_{1}$. Likewise, we find $v_{Z}$ such that $f^{X\top}v_{Z}^{\top}=1$ and $F^{X\top}v_{Z}^{\top}=\hat{e}_{i_{Z}}^{\top}$. It is then easy to verify that $(v_{X},v_{Z},v_{B})$ solves Eq. (36). ∎
@@
Ideally, one would also be able to show that the $Y$-system has good fault-distance if the individual $X$ and $Z$ systems do. However, the proof of fault-distance is complicated by the non-CSS nature of the merged code, meaning we lack a lemma similar to Lemma 10. In specific cases, one can check the fault distance by applying Lemma 9 with $\mathcal{U}=\langle\mathcal{S}_{O},\mathcal{S}_{Y}\rangle$ and $\mathcal{L}=\langle\bar{X}_{i},\bar{Z}_{i},i=1,2,\dots,k-1\rangle$, where $\mathcal{S}_{O}$ is the stabilizer group of the original code (with the additional qubits starting in $\ket{+}$ for the bottom half of Figure 7 and the bridge and in |$\ket{0}$ for the top half), $\mathcal{S}_{Y}$ the stabilizer group of the merged code shown in Figure 7, and $\mathcal{L}$ a simultaneous $k-1$ qubit logical basis for each of the codes $\mathcal{S}_{O}^{\prime}=\langle\mathcal{S}_{O},\bar{Y}_{M}\rangle$, $\mathcal{S}_{Y}$ and $\mathcal{U}$.
@@
We can at least prove that the code distance of the merged code with stabilizer group $\mathcal{S}_{Y}$ is good.

### v3

@@ 3.3 Code Distance and Fault Distance
3.4 Redundancy of Gauge Checks
As discussed in section 3.1, without further assumption on the induced Tanner graph $F$, we cannot prove that $\text{null}(F)$ has a low weight basis. This means we do not have a worst case upper bound on the weight of gauge checks in $U_{L}$. We believe that this is not an issue in average case applications, for two reasons. First, many of these gauge checks may be redundant and therefore do not need to be explicitly measured. Second, for a particularly heavy gauge check, we may decompose it into small gauge checks by a process we call cellulation.
We first discuss redundancy. Let $c\in C_{1}$ be such that $cF=0$, then $Z(c)$ is a gauge-fixed operator in the ancilla system. Observe that
$Z(c)\cdot\mathcal{H}_{Z}(c\in C_{0})=H_{Z}(c\in C_{0}),$
where the right hand side is a stabilizer in the original code $G$, with no support on $V_{0}$. If there exists a collection of checks $\bar{c}\subset\mathcal{C}_{0}^{Z}$ such that $H_{Z}(\bar{c})=H_{Z}(c)$, then the gauge check $Z(c)$ will be redundant in the merged code. To summarize, we have just shown the following lemma.
Lemma 12.
For a vector $c\in C_{0}$ with $cF=0$, if there exists vector $\bar{c}\in\mathcal{C}_{0}^{Z}$ such that $H_{Z}(\bar{c})=H_{Z}(c)$, then the gauge check $Z(c\in C_{1})$ is redundant in the merged code $\mathcal{G}_{X}$ and can be omitted from the Tanner graph.
This observation suggests an important heuristic for applying our scheme: if the stabilizers of the original code has large degrees of redundancy, then many (if not all) the gauge checks in our ancilla system may be omitted. We note that this redundancy condition is not hard to satisfy. Many important families of QLDPC codes, such as hypergraph product codes and lifted product codes, have structured redundancy among its stabilizer checks. We demonstrate this principle with the example of hypergraph product codes.
3.4.1 Hypergraph Product Codes
Consider a classical code specified by a binary $m\times n$ parity check matrix $H$. The hypergraph product code $Q=\text{HGP}(H,H^{\top})$ has parity check matrices $H_{Z}=[H\otimes I_{n},I_{m}\otimes H^{\top}]$ and $H_{X}=[I_{n}\otimes H,H^{\top}\otimes I_{m}]$. Suppose $\operatorname{rank}(H)=r$, then $Q$ is defined on $n^{2}+m^{2}$ physical qubits, and has logical dimension $(n-r)^{2}+(m-r)^{2}$.
We specify a basis for the $X$-logical operators of $Q$. Let $\operatorname{im}(\cdot)$ denote the image of a linear operator, we first choose bases for four linear spaces. Let $W=\{w_{1},\cdots,w_{n-r}\}$ be an arbitrary basis of $\ker{H}$, $V=\{v_{1},\cdots,v_{m-r}\}$ be an arbitrary basis of $\ker{H^{\top}}$. For the vector space $\mathbb{F}_{2}^{n}/\operatorname{im}(H^{\top})$, we take the reduced row echelon form of $H$, and let $J\subset[n]$ denote the indices of columns with pivots. We can take $R=\{e_{i}:i\in[n]\setminus J\}$, which is a basis of $\mathbb{F}_{2}^{n}/\operatorname{im}(H^{\top})$ such that every vector has weight one. For $\mathbb{F}_{2}^{m}/\operatorname{im}(H)$, we pick $S=\{s_{1},\cdots,s_{m-r}\}$ similarly. Now consider the sets
$X_{n\times n}$ =$=\{(w\otimes r,0_{m\times m}):w\in W\subset\ker(H),r\in R\subset\mathbb{F}_{2}^{n}/\operatorname{im}(H^{\top})\},$ (28)
$X_{m\times m}$ =$=\{(0_{n\times n},v\otimes s):v\in V\subset\ker(H^{\top}),s\in S\subset\mathbb{F}_{2}^{m}/\operatorname{im}(H)\}.$ (29)
We cite as a fact that these two sets specify a complete basis of the $X$-logical operators of $Q$. Note that $|X_{n\times n}|=(n-r)^{2}$ and $|X_{m\times m}|=(m-r)^{2}$.
Theorem 13.
For a gauge-fixed ancilla system on an $X$-operator $\bar{X}\in X_{n\times n}\cup X_{m\times m}$, all gauge checks are redundant.
Proof.
Fix an $X$-operator $P$ supported on $(w\otimes r,0_{m\times m})\in X_{n\times n}$, we will prove this claim for the gauge-fixed ancilla system built on $P$. The case where $P$ is supported on a basis vector in $X_{m\times m}$ follows similarly.
Consider the induced Tanner graph of $P$, which consists of qubits $V_{0}=\operatorname{supp}(w\otimes r)$ and all $Z$-stabilizer checks that touch qubits in $V_{0}$. By construction of the hypergraph product code $Q$, we can label the $Z$-checks of $Q$ by $f_{j}\otimes e_{i}$ where $f_{j}$ ranges over standard basis vectors of $\mathbb{F}_{2}^{m}$, and $e_{i}$ ranges over standard basis vectors of $\mathbb{F}_{2}^{n}$. The support of such a $Z$-stabilizer is
$H_{Z}(f_{j}\otimes e_{i})=((H^{\top}f_{j})\otimes e_{i},f_{j}\otimes(He_{i})).$ (30)
Let $C=\{f_{j}\in\mathbb{F}_{2}^{m}:H^{\top}f_{j}\cap w\neq\varnothing\}$, then $C\otimes r=\{f\otimes r:f\in C\}$ is exactly the set of $Z$-checks of $Q$ that touches $V_{0}$. Let $F$ be the submatrix of $H$ induced by $C$ and $\operatorname{supp}(w)$, namely, $F:C\rightarrow\operatorname{supp}(w)$. Since $r$ is a standard basis vector, we have
$F=F\otimes r:C\otimes r\rightarrow\operatorname{supp}(w\otimes r)$ (31)
is the submatrix of $H_{Z}$ induced by the qubits $V_{0}$ and $Z$-checks $C_{0}=C\otimes r$. Therefore, in a $L$ layer gauge-fixed ancilla system build on $P$, we can identify every $C_{i}$ as a copy of $C$, and every $V_{i}$ as a copy $\operatorname{supp}(w)$.
The gauge $Z$-check of our system corresponds to vectors $c\in\text{null}(F)\subset C$, i.e., $cF=0$. Fix a gauge check $Z(c)$, and consider its equivalent representative supported on qubits in $C_{1}$. We will show that $Z(c)$ is a product of $Z$-stabilizers in $\mathcal{C}_{Z}$. Let $\mathcal{H}_{Z}$ denote the new $Z$-check matrix with the ancilla system. Consider the product of $Z$-checks $c\otimes r\in\mathcal{C}^{Z}$, which is
$\mathcal{H}_{Z}(c\otimes r\in\mathcal{C}^{Z})$ $=Z(c\in C_{1})\cdot H_{Z}(c\otimes r)=Z(c\in C_{1})\cdot Z((H^{\top}c)\otimes r,c\otimes(Hr)).$ (32)
Since $c\in C$ is in null space of $F$, and $F$ is the submatrix of $H$ induced by $C$ and $\operatorname{supp}(w)$, we must have $H^{\top}c=0$. Therefore,
$Z(c\in C_{1})\cdot\mathcal{H}_{Z}(c\otimes r\in\mathcal{C}^{Z})$ $=Z(0_{n\times n},c\otimes(Hr)).$ (33)
By construction, $r\in\mathbb{F}_{2}^{n}/\operatorname{im}(H^{\top})$. Therefore, there does not exist $\bar{c}\in\mathbb{F}_{2}^{m}$ such that $H^{\top}\bar{c}=r$. In other words, $r$ is not in the row space of $H$, so $r$ is not orthogonal to the codespace $\ker(H)$. This implies that there exists vector $\bar{v}\in\ker(H)$ such that $\bar{v}\cdot r=1$. Let $v=\bar{v}-r$, then $Hv=Hr$. Consider the product of $Z$-checks $c\otimes v$, we have
$\mathcal{H}_{Z}(c\otimes v)$ $=H_{Z}(c\otimes v)$ (34)
$=Z((H^{\top}c)\otimes v,c\otimes(Hv))$ (35)
$=Z(0_{n\times n},c\otimes(Hr)).$ (36)
$Z(c\in C_{1})$ $=\mathcal{H}_{Z}(c\otimes v)\cdot\mathcal{H}_{Z}(c\otimes r\in\mathcal{C}^{Z}).$ (37)
Note that the first equation is true because the $Z$-checks in $c\otimes v$ are kept unchanged by the ancilla system, which is only attached to the checks $C\otimes r$. This proves that the gauge check $Z(c\in C_{1})$ is a product of $Z$-stabilizers in $\mathcal{C}^{Z}$. ∎
As a corollary, for hypergraph product codes the merged code $\mathcal{G_{X}}$ has degree $\delta+1$ if the original code has degree $\delta$. We expect that similar analysis can be made for lifted product codes and potentially other families of QLDPC codes. We further remark that Lemma 12 extends to all the joint measurement systems which we discuss later.
In cases where there is a heavy, non-redundant gauge check, a further technique called cellulation [Has21, SGI+24] could be employed. For a gauge check $Z(c)$, enumerate the checks in $c$ as $c_{1},\cdots,c_{k}$. For a subset of indices $I\subset[k]$, let $\bar{c}_{I}$ denote the product of checks $\prod_{i\in I}\mathcal{H}^{Z}(c_{i})$. If there exists $I\subset[k]$ such that $\bar{c}_{I}$ is low weight (where the definition of low depends on the use case), we can add $\bar{c}_{I}$ as a redundant check into the original code, and build the ancilla system on the modified code. This procedure breaks $Z(c)$ into a product of two gauge checks: one with $Z$ acting on $c_{i}$ for $i\in I$ and $\bar{c}_{I}$, the other with $Z$ acting on $c_{j}$ for $j\in[k]\setminus I$ and $\bar{c}_{I}$. Each of these gauge checks now have lower weight than $Z(c)$. We remark that [WY24] utilized cellulation in conjunction of another technique called decongestion [FH21] to guarantee that the measurement ancilla is LDPC for any QLDPC code.
@@ 3.6 Joint $X$ Measurement
We note that the weight of this basis of gauge checks $U^{B}$ we specified depends on the logical operators being measured. In this work, we do not attempt to optimize these weights, and note that in the worst case they may not be LDPC. We discuss the problem of minimizing weights of $U^{B}$, and the progress made by [SJOY24], in Section 3.8. In our study on the gross code, we found a basis of $U^{B}$ with weight at most $5$.
With the help of these gauge checks, we can show that the merged code indeed has $k-1$ logical qubits.
@@
In the merged code $\mathcal{G}_{XX}$, denote the number of layers in the $\bar{X}_{1}$ ($\bar{X}_{2}$) subsystem as $L_{1}$ ($L_{2}$), and the boundary Cheeger constant of $F^{(1)^{\top}}$ ($F^{(2)^{\top}}$) as $\beta_{1}$ ($\beta_{2}$). If $\lceil L_{1}/2\rceil\geq 1/\beta_{1}$, $\lceil L_{2}/2\rceil\geq 1/\beta_{2}$, and we measure the merged code for $R\geq d$ rounds, then the joint measurement system connecting those ancilla systems via a bridge (see Figure 6) has measurement and logical fault distances at least $d$.
@@
Let $\mathcal{S}$ be the stabilizer group of the original code and each ancilla qubit, including bridge qubits, in |$\ket{0}$, and $\mathcal{S}_{XX}$ be the stabilizer group of the joint merged code, Figure 4. Also let $\mathcal{L}$ be the group of logical operators of the unmeasured $k-1$ qubit logical space and =$\mathcal{L}^{*}=\mathcal{L}\setminus\{I\}$. The versions of Lemmas 9 and 10 that apply to the joint code imply that the measurement fault distance is $\min(R,d_{Z}(\bar{Z}_{1}\mathcal{L}\mathcal{S}))$ and the logical fault distance is $\min(d_{Z}(\mathcal{L}^{*}\mathcal{S}),d_{X}(\mathcal{L}^{*}\mathcal{S}_{XX}))$. The original code $\mathcal{S}$ has code distance $d$, therefore $d_{Z}(\bar{Z}_{1}\mathcal{L}\mathcal{S})\geq d$ and $d_{Z}(\mathcal{L}^{*}\mathcal{S})\geq d$. To complete the proof, it remains only to show that the $X$-distance of the joint merged code is good, i.e. $d_{X}(\mathcal{L}^{*}\mathcal{S}_{XX})\geq d$.
Let $P=X(t\in\mathcal{V})$ be an $X$-type logical operator of $\mathcal{G}$ that is inequivalent to $\bar{X}_{1}$ and $\bar{X}_{2}$ (note that $\bar{X}_{1}\equiv\bar{X}_{2}$). Consider the product of $P$ with a collection of $X$ checks, $\{v_{i}\in V_{i}^{(1)}\}_{\text{odd }i\in[L_{1}]}$ and $\{u_{j}\in V_{j}^{(2)}\}_{\text{odd }j\in[L_{2}]}$. Suppose $t\cap V_{0}^{(1)}=s_{1}$ and $t\cap V_{0}^{(2)}=s_{2}$. By the Expansion Lemma (Lemma 4), the weight of the new operator $P^{\prime}$ is
$|P^{\prime}|$ $=|X(t)\cdot\prod_{\begin{subarray}{c}i=1\\ i\text{ odd}\end{subarray}}^{L_{1}}\mathcal{H}_{X}(v_{i})\cdot\prod_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L_{2}}\mathcal{H}_{X}(u_{j})|$ (44)
$\geq|t|-|s_{1}|-|s_{2}|+\min(|s_{1}|,|V_{0}^{(1)}|-|s_{1}|)+\min(|s_{2}|,|V_{0}^{(2)}|-|s_{2}|)+|v_{L_{1}}+u_{L_{2}}|.$ (45)
If $|s_{1}|\geq|V_{0}^{(1)}|-|s_{1}|$ or $|s_{2}|\geq|V_{0}^{(2)}|-|s_{2}|$, then by Lemma 5 we have $|P^{\prime}|\geq d$. Otherwise, we have
$|P^{\prime}|$ $\geq|t|-|s_{1}|-|s_{2}|+|V_{0}^{(1)}|-|s_{1}|+|V_{0}^{(2)}|-|s_{2}|$ (46)
$=|P\cdot\bar{X}_{1}\cdot\bar{X}_{2}|\geq d,$ (47)
since the code distance of the original code is at least $d$. Now suppose $P\equiv\bar{X}_{1}\equiv\bar{X}_{2}$, then $P\equiv X(B)$. We will show that the weight of $X(B)$ cannot be reduced below $d$ by stabilizers in $\mathcal{S}_{XX}$.
Consider the product of $X(B)$ with a collection of $X$ checks in the ancilla system, $\{v_{i}\in V_{i}^{(1)}\}_{\text{odd }i\in[L_{1}]}$ and $\{u_{j}\in V_{j}^{(2)}\}_{\text{odd }j\in[L_{2}]}$. We again apply the expansion lemma, while also accounting for the support on $V_{0}^{(1)}$ and $V_{0}^{(2)}$.
$|P^{\prime}|$ $=|X(B)\cdot\prod_{\begin{subarray}{c}i=1\\ i\text{ odd}\end{subarray}}^{L_{1}}\mathcal{H}_{X}(v_{i})\cdot\prod_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L_{2}}\mathcal{H}_{X}(u_{j})|$ (48)
$\geq|B+S^{(1)}v_{L_{1}}+S^{(2)}v_{L_{2}}|+|v_{1}+v_{m}|+|u_{1}+u_{m}|+|v_{m}+v_{L_{1}}|+|u_{m}+u_{L_{2}}|$
$+\min(|v_{m}|,|\bar{X}_{1}|-|v_{m}|)+\min(|u_{m}|,|\bar{X}_{2}|-|u_{m}|)+|v_{1}|+|u_{1}|.$ (49)
$P^{\prime}$ now has support $v_{1}$ on $V_{0}^{(1)}$ and $u_{1}$ on $V_{0}^{(2)}$. If we multiply $P^{\prime}$ by $X$-stabilizers in $\mathcal{S}$, then this support could be further reduced. For any operator $O$, let $\mathcal{R}_{\mathcal{S}}(O)$ denote the minimum weight of equivalent representitives of $O$ under multiplication by stabilizers in $\mathcal{S}$. Then we have
$\mathcal{R}_{\mathcal{S}}(P^{\prime})$ $\geq|B+S^{(1)}v_{L_{1}}+S^{(2)}u_{L_{2}}|+|v_{1}+v_{m}|+|u_{1}+u_{m}|+|v_{m}+v_{L_{1}}|+|u_{m}+u_{L_{2}}|$
$+\min(|v_{m}|,|\bar{X}_{1}|-|v_{m}|)+\min(|u_{m}|,|\bar{X}_{2}|-|u_{m}|)+\mathcal{R}_{S}(v_{1}u_{1}).$ (50)
We do a case analysis based on $|v_{m}|,|u_{m}|$.
  1. 1.
    If $|v_{m}|\leq|\bar{X}_{1}|-|v_{m}|$ and $|u_{m}|\leq|\bar{X}_{2}|-|u_{m}|$, we have
    $\mathcal{R}_{\mathcal{S}}(P^{\prime})$ $\geq|B+S^{(1)}v_{L_{1}}+S^{(2)}u_{L_{2}}|+|v_{m}+v_{L_{1}}|+|u_{m}+u_{L_{2}}|+|v_{m}|+|u_{m}|$ (51)
    $\geq|B+S^{(1)}v_{L_{1}}+S^{(2)}u_{L_{2}}|+|v_{L_{1}}|+|u_{L_{2}}|$ (52)
    $\geq|B|\geq d.$ (53)
  2. 2.
    If $|v_{m}|\geq|\bar{X}_{1}|-|v_{m}|$ and $|u_{m}|\leq|\bar{X}_{2}|-|u_{m}|$, we have
    $\mathcal{R}_{\mathcal{S}}(P^{\prime})$ $\geq|u_{1}+u_{m}|+|u_{m}|+|\bar{X}_{1}|-|v_{m}|+\mathcal{R}_{\mathcal{S}}(v_{1}u_{1})$ (54)
    $\geq|u_{1}|+|\bar{X}_{1}|-|v_{1}|+\mathcal{R}_{\mathcal{S}}(v_{1}u_{1}).$ (55)
    Note that $\mathcal{R}_{\mathcal{S}}(v_{1}u_{1})+|u_{1}|\geq\mathcal{R}_{\mathcal{S}}(v_{1})$. Therefore $\mathcal{R}_{\mathcal{S}}(P^{\prime})\geq|\bar{X}_{1}|-|v_{1}|+\mathcal{R}_{\mathcal{S}}(v_{1})\geq\mathcal{R}_{\mathcal{S}}(\bar{X}_{1})\geq d$.
  3. 3.
    The case $|v_{m}|\leq|\bar{X}_{1}|-|v_{m}|$ and $|u_{m}|\geq|\bar{X}_{2}|-|u_{m}|$ follows similarly.
  4. 4.
    If $|v_{m}|\geq|\bar{X}_{1}|-|v_{m}|$ and $|u_{m}|\geq|\bar{X}_{2}|-|u_{m}|$, we have
    $\mathcal{R}_{\mathcal{S}}(P^{\prime})$ $\geq|\bar{X}_{1}|-|v_{m}|+|\bar{X}_{2}|-|u_{m}|+\mathcal{R}_{\mathcal{S}}(v_{1}u_{1})$ (56)
    $\geq|\bar{X}_{1}|-|v_{1}|+|\bar{X}_{2}|-|u_{1}|+\mathcal{R}_{\mathcal{S}}(v_{1}u_{1})$ (57)
    $\geq\mathcal{R}_{\mathcal{S}}(\bar{X}_{1}\cdot\bar{X}_{2})\geq d.$ (58)
This proves our claim. ∎
3.7 Gauge-fixed $Y$ Ancilla System
@@
$G=\left(\begin{array}[]{ccc}f^{Z^{\top}}&f^{X^{\top}}&0\\ F^{Z^{\top}}&0&S^{Z}\\ 0&F^{X^{\top}}&S^{X}\end{array}\right).$ (59)
@@
We claim that solutions $(v_{X},v_{Z},0)$ with $v_{B}=0$ are the sum of solutions in $U^{X}$ and $U^{Z}$. Note that, for any $v_{X}$, applying $X(v_{X}\in C_{1}^{X})$ flips an even number of checks from $V_{1}^{Z}\cup q_{1}$ (see Lemma 8). Therefore, because $F^{Z^{\top}}v_{X}^{\top}=0$, meaning $v_{X}$ flips no checks in $V_{1}^{Z}$, $v_{X}$ also does not flip $q_{1}$ or $f^{Z^{\top}}v_{X}^{\top}=0$. This shows $(v_{X},0,0)$ is a solution to Eq. (60). Likewise, it can be argued $(0,v_{Z},0)$ is a solution.
@@
Using Lemma 8, we can find a vector $v_{X}$ such that $f^{Z^{\top}}v_{X}^{\top}=1$ and $F^{Z^{\top}}v_{X}^{\top}=\hat{e}_{i_{X}}^{\top}$ because $X(v_{X}\in C_{1}^{X})$ flips exactly two checks, $i_{X}$ and $q_{1}$. Likewise, we find $v_{Z}$ such that $f^{X^{\top}}v_{Z}^{\top}=1$ and $F^{X^{\top}}v_{Z}^{\top}=\hat{e}_{i_{Z}}^{\top}$. It is then easy to verify that $(v_{X},v_{Z},v_{B})$ solves Eq. (60). ∎
@@
To show that the $Y$-system has good fault-distance, we apply Lemma 9 with $\mathcal{U}=\langle\mathcal{S}_{O},\mathcal{S}_{Y}\rangle$ and $\mathcal{L}=\langle\bar{X}_{i},\bar{Z}_{i},i=1,2,\dots,k-1\rangle$, where $\mathcal{S}_{O}$ is the stabilizer group of the original code (with the additional qubits starting in $\ket{+}$ for the bottom half of Figure 5 and the bridge and in |$\ket{0}$ for the top half), $\mathcal{S}_{Y}$ the stabilizer group of the merged code shown in Figure 5, and $\mathcal{L}$ a simultaneous $k-1$ qubit logical basis for each of the codes $\mathcal{S}_{O}^{\prime}=\langle\mathcal{S}_{O},\bar{Y}_{M}\rangle$, $\mathcal{S}_{Y}$ and $\mathcal{U}$. We have that the logical fault distance is at least $d(\mathcal{L}^{*}\mathcal{U})$ and the measurement fault distance is $\min(R,d(\bar{Z}_{M}\mathcal{L}\mathcal{U}),d(\bar{X}_{M}\mathcal{L}\mathcal{U}))$ for $R$ rounds of measurement.
Theorem 21.
If the $X$-system and $Z$-system individually have code distances at least $d$, then the $Y$-system has logical fault distance at least $d$.
Proof.
Given how the qubits in the ancilla system are initialized, let
$\operatorname{Init}=\langle Z(v)\text{ for singleton }v\in C^{Z}_{j\text{ odd}},X(v)\text{ for singleton }v\in C^{X}_{j\text{ odd}},X(b)\text{ for singleton }b\in B\rangle,$ (61)
where we used $C^{Z}_{j\text{ odd}}$ to mean $\bigcup_{j\text{ odd}}C^{Z}_{j}$ and similarly for the other variants of this notation. Then we have
$\mathcal{U}=\langle\mathcal{S}_{Y},\operatorname{Init}\rangle.$ (62)
Suppose $\bar{P}$ is a nontrivial logical operator of $\mathcal{G}_{Y}$, the merged code of the $Y$ system. Let $\bar{P}=P_{X}P_{Z}$ be its decomposition into an $X$-type Pauli $P_{X}$ and $Z$-type Pauli $P_{Z}$. Throughout we let $\mathcal{G}_{X}$ and $\mathcal{G}_{Z}$ denote the merged codes of the $X$- and $Z$-systems, respectively. When we write $P^{\prime}_{X}=P_{X}|_{\mathcal{G}_{X}}$ or $P^{\prime}_{Z}=P_{Z}|_{\mathcal{G}_{Z}}$, we mean to restrict those operators originally supported on $\mathcal{G}_{Y}$ to the qubits that exist in $\mathcal{G}_{X}$ or $\mathcal{G}_{Z}$, respectively. To show $d(\mathcal{L}^{*}\mathcal{U})\geq d$, we will show that either $|P^{\prime}_{X}|\geq d$ or $|P^{\prime}_{Z}|\geq d$. This is sufficient since $P^{\prime}_{X},P^{\prime}_{Z}$ are unaffected by operators in $\operatorname{Init}$, which means if $|P^{\prime}_{X}|\geq d$, then
$d(\mathcal{L}^{*}\mathcal{U})=\min_{\bar{P}\in\mathcal{L}^{*}\mathcal{S}_{Y}}\mathcal{R}_{\operatorname{Init}}(\bar{P})\geq\min_{\bar{P}\in\mathcal{L}^{*}\mathcal{S}_{Y}}\mathcal{R}_{\operatorname{Init}}(P^{\prime}_{X})\geq|P^{\prime}_{X}|\geq d.$ (63)
The first inequality is true since $\operatorname{Init}$ is generated by singleton checks.
We note that $P^{\prime}_{X}$ is a logical operator on $\mathcal{G}_{X}$. It commutes with all $Z$ checks of $\mathcal{G}_{X}$ because $\bar{P}$ commutes with all the checks of $\mathcal{G}_{Y}$. Similarly, $P^{\prime}_{Z}$ is a logical operator of $\mathcal{G}_{Z}$. We next claim that either $P^{\prime}_{X}$ is a nontrivial logical operator of $\mathcal{G}_{X}$ or $P^{\prime}_{Z}$ is a nontrivial logical operator of $\mathcal{G}_{Z}$. This can be seen by contradiction. Assuming both $P^{\prime}_{X}$ and $P^{\prime}_{Z}$ are products of stabilizers in their respective codes, then $\bar{P}$ can also be multiplied by the respective stabilizers from $\mathcal{G}_{Y}$ to find an equivalent operator with support that is much restricted:
Furthermore, by multiplying with gauge checks $U^{X},U^{Z},U^{B}$ and checks from $C^{Z}_{j\text{ even}}$ and $C^{X}_{i\text{ even}}$ we can move the support of $\bar{P}\equiv_{\mathcal{S}_{Y}}Z(b^{\prime}_{Z}\in B)X(b^{\prime}_{X}\in B)$ entirely onto the bridge qubits $B$ (this goes like the proof of Theorem 2 so we omit the details). However, in order to commute with all checks in $U^{B}$ (see Lemma 18), $V_{1}^{Z}$, and $V_{1}^{X}$, it must be that $b^{\prime}_{Z}=b^{\prime}_{X}=0$ or $\bar{P}\equiv_{\mathcal{S}_{Y}}I$. This is impossible because then $\bar{P}$ would not be a nontrivial logical operator of $\mathcal{G}_{Y}$.
Now to conclude the proof, we simply note that $P^{\prime}_{X}$ being nontrivial in $\mathcal{G}_{X}$ implies $|P^{\prime}_{X}|\geq d$ by assumption. Likewise, if $P^{\prime}_{Z}$ is nontrivial in $\mathcal{G}_{Z}$, we find $|P^{\prime}_{Z}|\geq d$. This completes our proof. ∎
We defer the proof of the following theorem to Appendix A.2, as it is similar to the proof of Theorem 17.
Theorem 22.
In the merged code $\mathcal{G}_{Y}$, denote the number of layers in the $\bar{X}$ ($\bar{Z}$) subsystem as $L_{X}$ ($L_{Z}$), and the boundary Cheeger constant of $(F^{X}f^{X})^{\top}$ ($(F^{Z}f^{Z})^{\top}$) as $\beta_{X}$ ($\beta_{Z}$). If $\lceil L_{X}/2\rceil\geq 1/\beta_{X}$, $\lceil L_{Z}/2\rceil\geq 1/\beta_{Z}$, and we measure the merged code for $R\geq d$ rounds, then the joint measurement system connecting those ancilla systems via a bridge (see Figure 5) has measurement fault distance at least $d$.
Remark 23.
@@
3.8 Bridging Codes
In this section, we consider joint measurements of logical operators on different code blocks. A version of our fault-tolerance guarantee is as follows. We defer the proof of this theorem to Appendix A.3, as it is similar to our earlier joint system proofs.